OpenStudy (anonymous):

vertex is (8,-3) also passes through (9,1) whats that in vertex form?

OpenStudy (anonymous):

you can use vertex form y=a(x-h)+k (h,k) is the vertex you can plug in a point for (x,y) and solve for a, then you'll have all parts of the equation =D so plugging in 1=a(9-8)-3 can you solve for a?

OpenStudy (anonymous):

1=1a-3?

OpenStudy (anonymous):

oopsy, i forgot the square, it'll still give us the same result in this case but want you to learn it right. its y=a(x-h)^2+k so we get 1=a(9-8)^2-3 1=1a-3 4=a just tack that into the equation a long with the vertex and you've got your answer y=4(x-8)^2-3

OpenStudy (anonymous):

So if im right that's the answer?

OpenStudy (anonymous):

yes that is the answer.

OpenStudy (anonymous):

thanks do you have time for another question?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

Your axis is x=6 and your maximum height is 8 and the line passes through 7,5

OpenStudy (anonymous):

I need standard form

OpenStudy (anonymous):

is that what the question says

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

my bad vertex form!

OpenStudy (anonymous):

axis of symmetry x=6?

OpenStudy (anonymous):

Are ya still here?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok well if I'm interpreting the question right, its an axis of symmetry. the problem basically gave you a vertex and a point like the last one because the axis of symmetry is the x value of the vertex, and the maximum value is the y value of the vertex. so we have a vertex of (6,8) and a point at (7,5) now its just like the last problem but with different numbers, can you start it off?

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

5=a(7-6)^2+8 solve for a.

OpenStudy (anonymous):

7/